Our home construction company still buys nails in 15-pound boxes but now we use an average of 3360 boxes a year. Preparing an order and receiving a shipment of nails involves a cost of $2.10 per order. Annual carrying costs are $0.6 per bag.

What will be their total cost of ordering and carrying the nails? (Include the pennies in your answer)
Answer #1

holding cost = average inventory level * holding cost per year

= 3360/2 * 0.6

= 1008

Economic order quantity = sqrt( 2*D*S/H)

= sqrt( 2 * 3360 * 2.1/1008)

= 3.74 = 4

total cost of ordering and carrying nails = Q/2 * H + D/Q * S

= 4/2 * 1008 + 3360/4 * 2.1

= 3780